OpenStudy (jacky90):

HELP Will medal and fan! In an electrical circuit, the current passing through a conductor varies inversely with the resistance. Suppose that when the current is 5 A (amperes), the resistance is 46 ohms. What is the resistance when the current is 2 A?

Hayhayz (hayhayz):

I=k/R

Hayhayz (hayhayz):

5=k/46 k=46*5=230 I=230/R 2=230/R 2*R=230 R=115

OpenStudy (jacky90):

TY! Love people who give the fast and easy answers!

Hayhayz (hayhayz):

yw :)